# CoSy: Evaluating Textual Explanations of Neurons

### v1

@@ 4.2 Do Models Respond Differently to Synthetic and Natural Images?
Given the visual similarity between natural and synthetic images of the same concepts, we investigate whether CV models respond differently to these groups and if the activation differences indicate adversarial behavior. To this end, we employed four different models pre-trained on ImageNet: ResNet18 [he2016deep], DenseNet161 [huang2017densely], GoogleNet [szegedy2015going], and ViT-B/16 [dosovitskiy2020image]. For each model, we randomly selected 10 output classes and generated 50 images per class using the class descriptions. We pass both synthetic and natural images through the models, collecting the activations of the output neuron corresponding to each class.
Figure 3: An overview of analyses performed to study the similarity between natural and synthetic images. From left to right: (1) an overview of MAD scores between synthetic and natural image activations of the output neuron’s ground truth classes for each model studied in this work, (2) activations collected for neuron 504 in ResNet18 for the class “coffee mug”, showcasing the difference between the natural and synthetic distributions and (3) examples of natural versus synthetic images. In both analyses, we observe a substantial overlap in the activations of synthetic and natural images, suggesting that the models respond similarly to both types of images.
Figure 3 (left) illustrates the distributions of the MAD between synthetic and natural images for the same class across the 10 classes. Generally, we observe that the activation of synthetic images is slightly higher than that of natural images of the same class. However, this difference is small, given the 0 value lies within 1 standard deviation. We also illustrated (Figure 3, right) the activations of neuron 504 in the ResNet18 output layer, corresponding to the “coffee mug” class. The results indicate a strong overlap in the neural response to both synthetic and natural images. While synthetic images activate the neuron slightly more, this doesn’t constitute an artifactual behavior or affect our framework, which we demonstrate in the following experiment.
4.3 Sanity Check
Table 2: Comparison of true and random explanations on output neurons with known ground truth labels. This table presents the quality estimates for true explanations, derived from target class labels, and random explanations, derived from randomly selected synthetic image classes (excluding the target class), across four models pre-trained on ImageNet. Higher values are better. Our results consistently show high scores for true explanations and low scores for random ones.
@@ 4.3 Sanity Check
ResNet18 0.94$\pm$0.20 0.48$\pm$0.23 15.95$\pm$6.88 -0.53$\pm$2.00
DenseNet161 0.95$\pm$0.17 0.52$\pm$0.24 15.27$\pm$5.88 -0.13$\pm$1.55
GoogLeNet 0.95$\pm$0.16 0.40$\pm$0.19 8.98$\pm$3.48 -0.41$\pm$0.57
ViT-B/16 0.97$\pm$0.11 0.52$\pm$0.16 7.55$\pm$2.69 -0.00$\pm$0.21
Following the experimental setup in Section 4.2, we selected a set of 10 output neurons and compared the CoSy scores of the ground truth explanations, given by the neuron label, with those of randomly selected explanations. The results, presented in Table 2, consistently demonstrate high scores for true explanations and low scores for random explanations. This experiment provides further evidence supporting the correctness of the proposed evaluation procedure. Additional experiments, including an analysis of the robustness of the evaluation measure, can be found in Appendix A.6.
Evaluating Explanation Methods
@@ 5.1 Benchmarking Explanation Methods
In this section, we evaluated three recent textual explanation methods, namely MILAN, INVERT, and CLIP-Dissect. Our analysis involves four distinct models: two pre-trained on the ImageNet dataset [russakovsky2015imagenet] (ResNet18 [he2016deep], ViT-B/16 [dosovitskiy2020image]) and two pre-trained on the Places365 dataset [zhou2017places] (DenseNet161 [huang2017densely], ResNet50 [he2016deep]). The ImageNet dataset focuses on objects, whereas the Places365 dataset is designed for scene recognition. Consequently, we customized our prompts accordingly: Prompt 5 performs best for object recognition, while for scene recognition, we found that Prompt 4 is more effective. Therefore, Prompt 4 was utilized in the Places365 experiment. For generating explanations with the explanation methods, we use a subset of 50,000 images from the training dataset on which the models were trained. For evaluation with CoSy, we use the corresponding validation datasets the models were pre-trained on as the control dataset. Additionally, for CLIP-Dissect, we define concept labels as a combination of the 20,000 most common English words and the corresponding dataset labels. For more details on Compute Resources refer to Appendix A.3.
Table 3: Benchmarking of explanation methods, explaining neurons on the second to last layers for different models. Explanations are generated with respect to a randomly selected set of 50 neurons where both AUC and MAD are reported. Higher values are better.
@@
ImageNet ResNet18 Avgpool MILAN 0.61$\pm$0.23 0.62$\pm$1.23
INVERT 0.93$\pm$0.11 2.88$\pm$1.54
CLIP-Dissect 0.93$\pm$0.11 3.45$\pm$1.71
ViT-B/16 Features MILAN 0.53$\pm$0.19 0.08$\pm$0.45
INVERT 0.89$\pm$0.17 0.97$\pm$0.46
CLIP-Dissect 0.78$\pm$0.19 0.76$\pm$0.62
Places365 DenseNet161 Features MILAN 0.56$\pm$0.28 0.08$\pm$0.27
INVERT 0.85$\pm$0.16 0.39$\pm$0.94
CLIP-Dissect 0.82$\pm$0.21 0.41$\pm$1.04
ResNet50 Avgpool MILAN 0.65$\pm$0.28 0.35$\pm$0.53
INVERT 0.94$\pm$0.08 1.04$\pm$0.74
CLIP-Dissect 0.92$\pm$0.11 1.05$\pm$0.75
Results of the evaluation can be found in Table 3. Overall, INVERT achieves the highest AUC scores across all models and datasets, except for the ResNet18 applied to ImageNet where CLIP-Dissect achieves a similar score. Also across other models and datasets, CLIP-Dissect demonstrates consistently good results. Since INVERT optimizes AUC in explanation generation, it may be biased towards AUC in our evaluation, leading to higher scores. MILAN generally performs poorly, with an average AUC below 0.65 across all tasks, indicating performance close to random guessing. This is somewhat expected since MILAN works with convolutional neurons. MILAN tends to generate highly abstract explanations, such as “white areas”, “nothing” or “similar patterns”. These abstract concepts are particularly challenging for a text-to-image model to generate accurately, likely contributing significantly to the low scores of MILAN. Contrary to the AUC scores, the MAD scores suggest that CLIP-Dissect outperforms INVERT for convolutional neural networks applied to both datasets. Nonetheless, in these cases, INVERT concepts also achieve consistently high scores. Otherwise, we find similar outcomes for both metrics $\Psi$, with MILAN achieving poor scores in all experimental settings.
@@ 5.2 Explanation Methods Struggle to Explain Lower Layer Neurons
In addition to the general benchmarking, we aimed to study the quality of explanations for neurons in different layers of a model. Since it is well known that lower-layer neurons usually encode lower-level concepts [lecun2015deep], it is interesting to see whether explanation methods can capture the concepts these neurons detect. To investigate this, we examined the quality of explanations across layers 1 to 4 and the output layer of an ImageNet pre-trained ResNet18. In addition to three prior explanation methods, we included the FALCON method in our analysis. For more details on the implementation of FALCON and FALCON-original see Appendix A.1.4. For each layer, we randomly selected 50 neurons for analysis.
Figure 4: A comparison of how different explanation methods vary in their quality, as measured by AUC (left) and MAD (right), across different layers in ResNet18. INVERT and CLIP-Dissect similarly high AUC and MAD scores across all layers, while MILAN, FALCON, and FALCON-original scores are comparably low. Generally, all methods perform increasingly worse on lower layers.
In Figure 4 we present the AUC (left) and MAD (right) results for all explanation methods across layers $1$ to $4$ and the output layer of ResNet18. While less pronounced for the AUC metric, in general, we find increasing scores for later layers across all methods and both metrics $\Psi$, which suggest higher concept quality in later layers. Furthermore, we find that similar to the benchmarking experiments, MILAN achieves lower scores across metrics. However, here, FALCON scores the lowest, not even surpassing random performance indicated by its AUC results (AUC$<0.5$).
@@ 5.3 What are Good Explanations?
Figure 5: A qualitative example, of neuron explanations across $4$ neurons. The first four panels include the textual explanation across INVERT, FALCON, CLIP-Dissect, and MILAN alongside three corresponding generated images. The respective AUC and MAD scores are reported below each panel. The last panel shows the activation distributions across $50$ generated images for each method and the distribution of the control data.
In our approach, we assume that testing visual representations of textual explanations on neurons can provide insights into what constitutes good explanations. Based on this assumption, we observe consistently high results from CLIP-Dissect and INVERT. The qualitative examples in Figure 5 demonstrate that their explanations share visually similar concepts (neurons $155$ and $459$) or even identical concepts (neuron $221$) while both achieving high AUC and MAD scores. It is important to note that although INVERT performs slightly better in several tasks, its applicability is limited to input data labels. In contrast, CLIP-Dissect can generate labels from a broader selection of concepts, though its reliance on a black-box model reduces interpretability compared to INVERT.
There are instances, such as neuron $260$ in Figure 5, where all explanations vary significantly. In these cases, we find that the explanation activation distributions of FALCON and MILAN often overlap with or even match the control dataset, providing the user with nearly random explanations. This observation aligns with our overall findings: both the AUC and MAD scores consistently indicate the low performance of FALCON and MILAN explanations in CoSy evaluation. Also, neurons $459$ and $155$ demonstrate the gap between consistently higher and lower-performing explanation methods, with AUC scores of 0.5 and below suggesting that these explanations are essentially random guesses.
Conclusion
In this work, we propose the first automatic evaluation framework for concept-based textual explanations of neurons. Unlike existing ad-hoc evaluation methods, we can now quantitatively compare different concept-based textual explanation methods against each other and test, whether the given explanation describes the neuron accurately, based on the neurons’s activations. We can evaluate the quality of individual neuron explanations by examining how accurately they align with the generated concept data points, without requiring human involvement.
Our comprehensive meta-evaluation demonstrates that CoSy guarantees a reliable explanation evaluation. In several experiments, we show that concept-based textual explanation methods are most applicable for the last layers, where high-level concepts are learned. In these layers, INVERT and CLIP-Dissect provide high-quality neuron concepts, whereas MILAN and FALCON explanations have lower quality and can present close to random concepts, which might lead to wrong conclusions about the network. Thus, the results highlight the importance of evaluation when using concept-based textual explanation methods.
@@ Conclusion
While we can present promising results, one of the key limitations of CoSy is the generative model. For example, the text-to-image model training might not include the generated concepts. This absence leads to worsened generative performance but could be circumvented by an analysis of pre-training datasets and model performance. Moreover, the model’s capabilities of generating highly abstract concepts like “white objects” are limited. In both cases, exploring more sophisticated, specialized, or constrained models could help.
@@
Evaluation of non-local explanation methods is still a largely neglected research area, where CoSy plays an important yet preliminary part. In the future, we need additional, complementary definitions of explanation quality that extend our precise definition of AUC and MAD, e.g., that involve humans to assess plausibility [Cheng2023] or evaluate explanation quality via the success of a downstream task [krishna2023post]. Furthermore, we plan to extend the application of our evaluation framework to additional domains including NLP and healthcare. In particular, it would be interesting to analyze the quality of more recent autointerpretable explanation methods given by highly opaque, large language models (LLMs) [Kroeger23, bills2023language]. Also, we believe that applying CoSy to healthcare datasets, where the quality of the explanation really matter, is an impactful next step.
@@ Appendix A Appendix
A.1 Concept-based Textual Explanation Methods
Concept-based textual explanation methods aim to provide insights into human-understandable concepts learned by DNNs, enabling a deeper understanding of their decision-making mechanisms. These methods provide textual descriptions for neurons in CV models. This creates a connection between the abstract representation of a concept by the neural network and a human interpretation. In general, a concept can be any abstraction, such as a color, an object, or even an idea [molnar2022]. Concept-based textual descriptions of a neuron $f_{i}$ can originate from various spaces depending on their generation process.
@@ A.1 Concept-based Textual Explanation Methods
Examples for such explanation methods are MILAN [hernandez2022natural], FALCON [pmlr-v202-kalibhat23a], CLIP-Dissect [oikarinen2023clip], and INVERT [bykov2023labeling]. Figure 6 shows the general principle of how $\mathcal{E}$ works. In Table 4 we outline the origin of textual descriptions and their corresponding set memberships for each $\mathcal{E}$.
Figure 6: Concept-Based Textuaĺ Explanation Methods. A neural representation $f_{i}$ is selected, and a concept-based textual explanation method $\mathcal{E}$ is applied to generate a textual description $s$ explaining $f_{i}$.
Table 4: Set Membership and Origin of Descriptions. Generated textual descriptions $s$ have varying set membership and origin across all $\mathcal{E}$. These descriptions can originate from distinct spaces: individual concepts $\mathcal{C}$, logical combinations of concepts $\mathcal{L}$, and open-ended natural language concept descriptions $\mathcal{N}$. A labeled dataset refers to a collection of images paired with individual concept labels. Generated captions are produced by image-to-text models, such as Show-Attend-Tell [xu2015show]. An image caption dataset consists of image-caption pairs. A concept set consists of textual concept labels.

### v2

@@ 4.2 Do Models Respond Differently to Synthetic and Natural Images?
Given the visual similarity between natural and synthetic images of the same class, we investigate whether CV models respond differently to these groups and if the activation differences indicate adversarial behavior. To this end, we employed four different models pre-trained on ImageNet: ResNet18 [37], DenseNet161 [38], GoogleNet [39], and ViT-B/16 [40]. For each model, we randomly selected 50 output classes and generated 50 images per class using the class descriptions. We pass both synthetic and natural images through the models, collecting the activations of the output neuron corresponding to each class.
Figure 3: An overview of analyses performed to study the similarity between natural and synthetic images. From left to right: (a) an overview of MAD scores between synthetic and natural image activations of the output neuron’s ground truth classes for each model studied in this work, (b) activations collected for neuron $504$ in ResNet18 for the class “coffee mug”, showcasing the difference between the natural and synthetic distributions and (c) examples of natural versus synthetic images. In both analyses, we observe a substantial overlap in the activations of synthetic and natural images, suggesting that the models respond similarly to both types of images.
Figure 3 (a) illustrates the distributions of the MAD between synthetic and natural images for the same class across the 50 classes. Across all models, we observe that the median activation of synthetic images is slightly higher than that of natural images of the same class. However, this difference is small, given the 0 value lies within 1 standard deviation. We also illustrate the activations of neuron $504$ in the ResNet18 output layer for the “coffee mug” class in Figure 3 (b). The results indicate a strong overlap in the neural response to both synthetic and natural images. While synthetic images activate the neuron slightly more, this does not constitute an artifactual behavior or affect our framework, which we demonstrate in the following experiment.
4.3 Random Baseline
Following the experimental setup in Section 4.2, we selected a set of 50 output neurons and compared the CoSy scores of the ground truth explanations, given by the neuron label, with those of randomly selected explanations. The results, presented in Table 2, consistently demonstrate high scores for true explanations and low scores for random explanations. This experiment provides further evidence supporting the correctness of the proposed evaluation procedure. An additional experiment that excludes the target class from the control dataset is presented in Appendix A.4, along with an analysis of the robustness of the evaluation measure detailed in Appendix A.6.
Table 2: Comparison of true and random explanations on output neurons with known ground truth labels. This table presents the average quality scores (with standard deviations) for true explanations, derived from target class labels, and random explanations, derived from randomly selected synthetic image classes (including the target class), across four models pre-trained on ImageNet. Higher values are better. Our results consistently show high scores for true explanations and low scores for random ones.
@@ 4.3 Random Baseline
ResNet18 0.98$\pm$0.09 0.47$\pm$0.21 6.46$\pm$2.07 -0.11$\pm$0.79
DenseNet161 0.99$\pm$0.08 0.44$\pm$0.22 7.11$\pm$1.82 -0.19$\pm$0.73
GoogLeNet 0.99$\pm$0.07 0.48$\pm$0.23 7.74$\pm$2.14 -0.06$\pm$0.78
ViT-B/16 0.99$\pm$0.05 0.49$\pm$0.22 13.12$\pm$3.19 0.09$\pm$1.10
5 Evaluating Explanation Methods
@@ 5.1 Benchmarking Explanation Methods
In this section, we evaluated three recent textual explanation methods, namely MILAN, INVERT, and CLIP-Dissect. Our analysis involves six distinct models: four pre-trained on the ImageNet dataset [36] (ResNet18 [37], ResNet50 [37], ViT-B/16 [40], DINO ViT-S/8 [41]) and two pre-trained on the Places365 dataset [42] (DenseNet161 [38], ResNet50 [37]). The ImageNet dataset focuses on objects, whereas the Places365 dataset is designed for scene recognition. Consequently, we customized our prompts accordingly: Prompt 5 performs best for object recognition, while for scene recognition, we found that Prompt 3 is more effective. Therefore, Prompt 3 was utilized in the Places365 experiment.
For generating explanations with the explanation methods, we use a subset of 50,000 images from the training dataset on which the models were trained. For evaluation with CoSy, we use the corresponding validation datasets the models were pre-trained on as the control dataset. Additionally, for CLIP-Dissect, we define concept labels by combining the 20,000 most common English words with the corresponding dataset labels. For INVERT we set the compositional length of the explanation as $L=1$, where $L\in\mathbb{N}$. For more details on compute resources, refer to Appendix A.7.
Table 3: Benchmarking of neuron description methods, for neurons in the second to last layer across different models. Explanations are generated for a randomly selected set of 50 neurons, with average scores for both AUC and MAD reported alongside standard deviations. Higher values indicate better performance; bold numbers represent the highest scores.
@@
ImageNet ResNet18 Avgpool MILAN 0.61$\pm$0.23 0.69$\pm$1.35
CLIP-Dissect 0.93$\pm$0.11 3.85$\pm$1.88
INVERT 0.93$\pm$0.11 3.23$\pm$1.72
ResNet50 Avgpool MILAN 0.44$\pm$0.23 -0.08$\pm$0.72
CLIP-Dissect 0.95$\pm$0.08 4.98$\pm$2.57
INVERT 0.96$\pm$0.06 4.62$\pm$2.26
ViT-B/16 Features MILAN 0.53$\pm$0.19 0.12$\pm$0.76
CLIP-Dissect 0.78$\pm$0.19 1.29$\pm$1.01
INVERT 0.89$\pm$0.17 1.67$\pm$0.82
DINO ViT-S/8 Layer 11 MILAN 0.59$\pm$0.21 0.37$\pm$0.91
CLIP-Dissect 0.95$\pm$0.08 4.59$\pm$2.62
INVERT 0.73$\pm$0.27 2.70$\pm$3.48
Places365 DenseNet161 Features MILAN 0.56$\pm$0.28 0.44$\pm$1.30
CLIP-Dissect 0.82$\pm$0.21 2.52$\pm$2.33
INVERT 0.85$\pm$0.16 2.21$\pm$1.95
ResNet50 Avgpool MILAN 0.65$\pm$0.28 1.11$\pm$1.67
CLIP-Dissect 0.92$\pm$0.11 3.73$\pm$2.39
INVERT 0.94$\pm$0.08 3.54$\pm$1.99
Results of the evaluation can be found in Table 3. Overall, INVERT achieves the highest AUC scores across all models and datasets, except for DINO ViT-S/8 and ResNet18 applied to ImageNet, where CLIP-Dissect achieves a higher or similar score. Also across other models and datasets, CLIP-Dissect demonstrates consistently good results. Since INVERT optimizes AUC in explanation generation, it may be biased towards AUC in our evaluation, leading to higher scores. MILAN generally performs poorly, with an average AUC below 0.65 across all tasks, indicating performance close to random guessing. MILAN tends to generate highly abstract explanations, such as “white areas”, “nothing” or “similar patterns”. These abstract concepts are particularly challenging for a text-to-image model to generate accurately, likely contributing significantly to the low scores of MILAN. Contrary to the AUC scores, the MAD scores suggest that CLIP-Dissect outperforms INVERT for convolutional neural networks applied to both datasets. Nonetheless, in these cases, INVERT concepts also achieve consistently high scores. Otherwise, we find similar outcomes for both metrics $\Psi$, with MILAN achieving poor scores in all experimental settings.
@@ 5.2 Explanation Methods Struggle to Explain Lower Layer Neurons
In addition to the general benchmarking, we aimed to study the quality of explanations for neurons in different layers of a model. Since it is well known that lower-layer neurons usually encode lower-level concepts [43], it is interesting to see whether explanation methods can capture the concepts these neurons detect. To investigate this, we examined the quality of explanations across layers 1 to 4 and the output layer of an ImageNet pre-trained ResNet18. In addition to three prior explanation methods, we included the FALCON method in our analysis. For more details on the implementation of FALCON see Appendix A.1.4, for additional results of the original FALCON implementation see Appendix A.8, and for qualitative examples and a discussion of lower-level concepts see Appendix A.9. For each layer, we randomly selected 50 neurons for analysis.
In Figure 4 we present the AUC and MAD results for all explanation methods across layers $1$ to $4$ and the output layer of ResNet18. While less pronounced for the AUC metric, in general, we find increasing scores for later layers across all methods and both metrics $\Psi$, which suggest higher concept quality in later layers. Furthermore, we find that similar to the benchmarking experiments, MILAN achieves lower scores across metrics. Both MILAN and FALCON consistently show lower performance, with AUC scores of $0.5$ indicating random guessing. Nonetheless, we point out that these methods typically output semantically high-level concepts. Potentially, this is related to the inherent difficulty in describing low-level abstractions in natural language given their complexity (see Appendix A.10).
Figure 4: A comparison of how different explanation methods vary in their quality, as measured by (a) AUC and (b) MAD, across different layers in ResNet18. INVERT and CLIP-Dissect maintain high AUC and MAD scores across all layers, while MILAN and FALCON have lower scores. Overall, performance declines in the lower layers for all methods.
@@ 5.3 What are Good Explanations?
Figure 5: A qualitative example, of neuron explanations across four neurons. The first four panels include the textual explanation across INVERT, FALCON, CLIP-Dissect, and MILAN alongside three corresponding generated images. The respective AUC and MAD scores are reported below each panel. The last panel shows the activation distributions across $50$ generated images for each method and the distribution of the control data.
In our approach, we propose that testing visual representations of textual explanations on neurons can provide insights into what constitutes good explanations. Building on this premise, we observe consistently high results from CLIP-Dissect and INVERT. The qualitative examples in Figure 5 demonstrate that their explanations share visually similar concepts (neurons $155$ and $459$) or even identical concepts (neuron $221$) while both achieving high AUC and MAD scores. It is important to note that although INVERT performs slightly better in several tasks, the explanations are constrained to the input data labels. In contrast, CLIP-Dissect can generate descriptions from a broader selection of concepts, though its reliance on a black-box model reduces interpretability compared to INVERT.
There are instances, such as neuron $260$ in Figure 5, where all explanations vary significantly. In these cases, we find that the explanation activation distributions of FALCON and MILAN often overlap with or even match the control dataset, providing the user with nearly random explanations. This observation aligns with our overall findings: both the AUC and MAD scores consistently reflect the low performance of FALCON and MILAN explanations in the CoSy evaluation. Also, neurons $459$ and $155$ demonstrate the gap between consistently higher and lower-performing explanation methods.
6 Conclusion
In this work, we propose the first automatic evaluation framework for textual explanations of neurons. Unlike existing ad-hoc evaluation methods, we can now quantitatively compare different neuron description methods against each other and test, whether the given explanation describes the neuron accurately, based on its activations. We can evaluate the quality of individual neuron explanations by examining how accurately they align with the generated concept data points, without requiring human involvement.
Our comprehensive sanity checks demonstrate that CoSy guarantees a reliable explanation evaluation. In several experiments, we show that neuron description methods are most applicable for the last layers, where high-level concepts are learned. In these layers, INVERT and CLIP-Dissect provide high-quality neuron concepts, whereas MILAN and FALCON explanations have lower quality and can present close to random concepts, which might lead to wrong conclusions about the network. Thus, the results highlight the importance of evaluation when using neuron description methods.
@@ 6 Conclusion
The use of generative models involves a distinct set of limitations. For instance, text-to-image models may not include certain concepts in their training data, which can reduce their generative performance. This limitation, however, can often be addressed by analyzing the pre-training datasets and assessing model performance. Moreover, the model’s capabilities of generating highly abstract concepts like “white objects” can be limited. However, the challenges with abstract concepts also reflect the descriptive quality of the provided explanations—explanations should be inherently understandable to humans. In both cases, exploring more sophisticated, specialized, or constrained models may offer improvement.
@@
Evaluation of non-local explanation methods is still a largely neglected research area, where CoSy plays an important yet preliminary part. In the future, we need additional, complementary definitions of explanation quality that extend our precise definition of AUC and MAD, e.g., that involve humans to assess plausibility [44] or evaluate explanation quality via the success of a downstream task [45]. Furthermore, we plan to extend the application of our evaluation framework to additional domains including NLP and healthcare. In particular, it would be interesting to analyze the quality of more recent autointerpretable explanation methods given by highly opaque, large language models (LLMs) [46, 9]. We believe that applying CoSy to healthcare datasets, where high-quality explanations are crucial, represents an impactful next step.
@@ Appendix A Appendix
A.1 Neuron Description Methods
Neuron description methods aim to provide insights into human-understandable concepts learned by DNNs, enabling a deeper understanding of their decision-making mechanisms. These methods provide textual descriptions for neurons in CV models. This creates a connection between the abstract representation of a concept by the neural network and a human interpretation. In general, a concept can be any abstraction, such as a color, an object, or even an idea [47]. Textual explanations of a neuron $f_{i}$ can originate from various spaces depending on their generation process.
@@ A.1 Neuron Description Methods
Examples for such explanation methods are MILAN [13], FALCON [14], CLIP-Dissect [15], and INVERT [16]. Figure 6 shows the general principle of how $\mathcal{E}$ works. In Table 4 we outline the origin of textual descriptions and their corresponding set memberships for each $\mathcal{E}$.
Figure 6: Neuron Description Methods. A neuron $f_{i}$ is selected, and a neuron description method $\mathcal{E}$ is applied to generate a textual description $s$ explaining $f_{i}$.
Table 4: Set Membership and Origin of Descriptions. Generated textual descriptions $s$ have varying set membership and origin across all $\mathcal{E}$. These descriptions can originate from distinct spaces: individual concepts $\mathcal{C}$, logical combinations of concepts $\mathcal{L}$, and open-ended natural language concept descriptions $\mathcal{N}$. A labeled dataset refers to a collection of images paired with individual concept labels. Generated captions are produced by image-to-text models, such as Show-Attend-Tell [48]. An image caption dataset consists of image-caption pairs. A concept set consists of textual concept labels.